A copper wire is 37.79 cm long and weighs 1.589 g. The density of copper is 8.933 g/cm3.

What is the radius of the wire in millimeters?

?=.   mm

What is the diameter of the wire?

?=. mm
Answer #1

1)

1st find the volume of wire:

volume = mass / density

= 1.589 g / (8.933 g/cm^3)

= 0.1779 cm^3

use:

volume = pi*r^2*l

0.1779 = 3.14 * r^2 * 37.79

r^2 = 1.499*10^-3

r = 0.03872 cm

r = 0.3872 mm

Answer: 0.3872 mm

2)

diameter = 2*radius

= 2*0.3872 mm

= 0.7744 mm

Answer: 0.7744 mm